Literaturverzeichnis 151

[89] M. V. Berry und J. P. Keating, J. Phys. A 23, 4839 (1990).

[90] J. P. Keating, Chaos 2, 15 (1992).

[91] J. P. Keating, Proc. R. Soc. Lond. A 436, 99 (1992).

[92] H.-J. Sommers, F. Haake und J. Weber, in Vorbereitung.

[93] M. V. Berry, Phil. Trans. R. Soc. 287, 237 (1977).

[94] M. Wilkinson, J. Phys. A 21, 1173 (1988).

[95] N. Argaman, F.-M. Dittes, E. Doron, J. P. Keating, A. Yu. Kitaev, M. Sieber und U. Smilansky, Phys. Rev. Lett. 71, 4326 (1993).

[96] A. V. Andreev und B. L. Altshuler, Phys. Rev. Lett. 75, 902 (1995).

[97] O. Agam, B. L. Altshuler und A. V. Andreev, Phys. Rev. Lett. 75, 4389 (1995).

[98] A. V. Andreev, O. Agam, B. D. Simons und B. L. Altshuler, Phys. Rev. Lett. 76, 3947 (1996).

[99] A. V. Andreev, B. D. Simons, O. Agam und B. L. Altshuler, cond-mat/9605204 (1996).

[100] E. Bogomolny und J. P. Keating, Phys. Rev. Lett. 77, 1472 (1996).

[101] U. Smilansky, chao-dyn/9611002 (1996).

[102] J. H. Hannay und A. M. Ozorio de Almeida, J. Phys. A 17, 3429 (1984).

[103] E. B. Bogomolny, Comments At. Mol. Phys. 25, 67 (1990).

[104] E. B. Bogomolny, Nonlinearity 5, 805 (1992).

[105] S. C. Creagh und R. G. Littlejohn, Phys. Rev. A 44, 836 (1991).

[106] S. C. Creagh und R. G. Littlejohn, J. Phys. A 25, 1643 (1992).